OpenStudy (anonymous):

I'm having issues with this trig integral... Integral of (x^2)dx/(sqrt(x^2 + 1))

OpenStudy (amistre64):

what have you tried?

OpenStudy (anonymous):

I've tried various things but I get stuck when I get it to look like:

OpenStudy (anonymous):

\[\int\limits \tan^2\theta \sec \theta d \theta\]

OpenStudy (anonymous):

I used \[\sec \theta = \sqrt{\tan^2\theta + 1}\] and used x = \[\tan \theta\] and dx = \[\sec ^2 \theta\ d\theta\]

OpenStudy (amistre64):

change tan^2 to sec^2 -1; maybe? sec(sec^2 -1) = sec^3 - sec, that might not be easier tho

OpenStudy (amistre64):

the derivative of sec = sec tan

OpenStudy (amistre64):

ugh .. but you got tan tan sec

myininaya (myininaya):